if you have 36 ft of fencing,what are the areas of different rectangles you could enclose with the fencing?
Answer by finly36(7)

total area=30
ok
a=l*w
so 2*3=6 ok 24 ft left
2*4=8 16ft left
4*4=16 rectangles
PS. there are many more ways to solve this this is just one way